# A Survey on In-context Learning

### v4

@@ A Survey on In-context Learning
Qingxiu Dong1, Lei Li1, Damai Dai1, Ce Zheng1, Jingyuan Ma1, Rui Li1, Heming Xia2, Jingjing Xu3, Zhiyong Wu4, Baobao Chang1, Xu Sun1, Lei Li5 and Zhifang Sui1 1 MOE Key Lab of Computational Linguistics, School of Computer Science, Peking University 2 The Hong Kong Polytechnic University 3 ByteDance 4 Shanghai AI Lab   5 Carnegie Mellon University dqx@stu.pku.edu.cn, szf@pku.edu.cn
@@ 1 Introduction
With the rapid growth of studies in ICL, our survey aims to sensitize the community toward the current progress. In the following sections, we delve into an in-depth discussion of related studies, and we summarize the key findings in Appendix A. We highlight the challenges and potential directions and hope our work provide a useful roadmap for beginners interested in this area and shed light on future research.
@@ 4.1.1 Demonstration Selection
Table 2: Fair comparison of demonstration selection methods. CQA and News are abbreviations of Commonsense QA and AG News, respectively. The best results are bolded. Our experiments on topk (Liu et al., 2022), votek (Su et al., 2023), mdl (Wu et al., 2023b) show that topk selects the most effective examples on average.
@@ 4.3 Scoring Function
Table 3: Summary of different scoring functions. Coverage refers to task coverage.
The scoring function determines how to transform the predictions of a language model into an estimation of the likelihood of a specific answer. The Direct method uses the conditional probability of candidate answers represented by tokens in the model’s vocabulary (Brown et al., 2020). The answer with the highest probability is selected as the final answer, but this method restricts template design by requiring answer tokens to be at the end of input sequences. Perplexity (PPL) is another commonly used metric that computes the sentence perplexity of the entire input sequence $S_{j}=\{C,s(x,y_{j},I)\}$, which includes tokens from demonstration examples $C$, the input query $x$, and the candidate label $y_{j}$. PPL evaluates the probability of the sentence, eliminating token position limitations but requiring additional computation time. Min et al. (2022a) proposed using channel models (Channel) to compute the conditional probability in reverse, estimating the likelihood of the input query given the label. This approach requires language models to generate every token in the input, potentially boosting performance under imbalanced training data. We summarize all three scoring functions in Table 3.
@@ 8 Conclusion
This paper offers a comprehensive examination and summary of current methodologies and analyses in the area of In-Context Learning (ICL). However, given the extensive body of related work, particularly in demonstration design and the principle analysis of ICL, we may have overlooked some equally valuable contributions. Additionally, we outline several future directions for research in ICL, including long-context ICL, efficiency and scalability in ICL, etc. We plan to leave these aspects for future work.

### v6

@@ A Survey on In-context Learning
Qingxiu Dong1, Lei Li1, Damai Dai1, Ce Zheng1, Jingyuan Ma1, Rui Li1, Heming Xia2, Jingjing Xu3, Zhiyong Wu4, Tianyu Liu5, Baobao Chang1, Xu Sun1, Lei Li6 and Zhifang Sui1 1 Peking University 2 The Hong Kong Polytechnic University 3 ByteDance 4 Shanghai AI Lab   5 Alibaba Group 6 Carnegie Mellon University dqx@stu.pku.edu.cn, szf@pku.edu.cn
11footnotetext: We list the author contributions and roles in Appendix D.4.
@@ 1 Introduction
With the rapid growth of studies in ICL, our survey aims to sensitize the community toward the current progress. In the following sections, we delve into an in-depth discussion of related studies, and we summarize the taxonomy in Figure 2 and the key findings in Appendix A. We highlight the challenges and potential directions and hope our work provide a useful roadmap for beginners interested in this area and shed light on future research.
@@ 4.1.1 Demonstration Selection
Table 2: Fair comparison of demonstration selection methods. CQA and News are abbreviations of Commonsense QA and AG News, respectively. The best results are bolded. Our experiments on topk (Liu et al., 2022), votek (Su et al., 2023), mdl (Wu et al., 2023b) show that the effectiveness of ICL example selection methods are model-dependent. On GPT-2, the mdl method performs the best, while on the other three models, topk performs the best.
@@ 4.3 Scoring Function
Table 3: Summary of different scoring functions. Coverage refers to task coverage. The qualitative results for ‘Efficiency’ and ‘Stability’ metrics are elaborated in Table 4 and Table 5, respectively.
The scoring function determines how to transform the predictions of a language model into an estimation of the likelihood of a specific answer. The Direct method uses the conditional probability of candidate answers represented by tokens in the model’s vocabulary (Brown et al., 2020). The answer with the highest probability is selected as the final answer, but this method restricts template design by requiring answer tokens to be at the end of input sequences. Perplexity (PPL) is another commonly used metric that computes the sentence perplexity of the entire input sequence $S_{j}=\{C,s(x,y_{j},I)\}$, which includes tokens from demonstration examples $C$, the input query $x$, and the candidate label $y_{j}$. PPL evaluates the probability of the sentence, eliminating token position limitations but requiring additional computation time. Min et al. (2022a) proposed using channel models (Channel) to compute the conditional probability in reverse, estimating the likelihood of the input query given the label. This approach requires language models to generate every token in the input, potentially boosting performance under imbalanced training data. We summarize all three scoring functions in Table 3. Note that in Table 3, ‘Efficiency’ refers to the language model inference latency; ‘Coverage’ reflects whether the method utilizes the output probability of the local or all token positions in the input sequence; and ‘Stability’ indicates whether the in-context learning ability is easily affected by changes in the demonstration examples.
@@ 8 Conclusion
This paper offers a comprehensive examination and summary of current methodologies and analyses in the area of In-Context Learning (ICL). However, given the extensive body of related work, particularly in demonstration design and the principle analysis of ICL, we may have overlooked some equally valuable contributions. Additionally, we outline several future directions for research in ICL, including long-context ICL, efficiency and scalability in ICL, etc. We plan to leave these aspects for future work. Furthermore, many papers covered by this survey did not utilize the most up-to-date models while running experiments. We advocate for more thorough and up-to-date research to provide actionable insights for practitioners.
@@ Appendix B Experimental Detail
Table 4 and Table 5 show the quantitative results on the efficiency and stability metrics for different scoring functions in Table 3.
Model Direct PPL Channel
GPT2 44.13(1.00) 114.02(2.58) 157.70(3.57)
GPT-J 611.04(1.00) 1766.82(2.89) 1793.27(2.93)
Qwen2 745.89(1.00) 1886.63(2.53) 1957.97(2.63)
Llama3 790.46(1.00) 1935.04(2.45) 1956.21(2.47)
AVG 1.00 2.61 2.90
Table 4: The qualitative results of the Efficiency metric in Table 3 which record the language model inference latency (including the time for scoring with different scoring functions, with input data containing 8 in-context examples). The unit is milliseconds (ms). Each cell’s parentheses contain the ratio of the latency for the current column model using the current row scoring function to the latency using direct inference. The final calculated average is the average of these ratios.
Model Direct PPL Channel
GPT2 1.12 0.85 3.18
GPT-J 1.00 0.77 4.06
Qwen2 0.72 0.70 2.43
Llama3 0.89 0.78 2.43
AVG 0.93 0.78 3.03
Table 5: The qualitative results of the Stability metric in Table 3 which reflect whether the in-context learning ability is easily affected by changes in demonstration examples. We conducted experiments using a test set of size 10k and set up 5 different random seeds. Each time, 8 examples were randomly selected from 5k training examples for the experiments. The table records the variance of performance.
@@ D.3 Speech In-Context Learning
D.4 Author Contributions
Qingxiu Dong led the project. Qingxiu Dong, Lei Li, Damai Dai, and Ce Zheng discussed and wrote the initial draft of the paper and the second version update. Jingyuan Ma, Rui Li, Heming Xia, and Qingxiu Dong discussed and wrote the third version update, incorporating new work that emerged over the year. Tianyu Liu was responsible for the fourth version update, adding new relevant work and polishing the manuscript. Jingjing Xu, Zhiyong Wu, and Lei Li (CMU) contributed to the revisions and discussions of the first two versions of the paper. Baobao Chang, Xu Sun, Lei Li (CMU), and Zhifang Sui guided the research of Qingxiu Dong, Lei Li, Damai Dai, and Ce Zheng.
D.5 Comparison with other survey papers
Our survey was drafted and posted on the Arxiv at the end of 2022, which is, to the best of our knowledge, the very first to review in-context learning in the field. We also regularly update this survey in a timely manner, with four major revisions.
Starting from 2023, we notice the emerge of several related survey in the field of in-context learning. Xu et al. (2024) made a comprehensive review on the choices for models, training procedures and inference algorithms to retrieve demonstrative examples of in-context learning. Li (2023) provided practical suggestions on prompt engineering for in-context learning. Zhou et al. (2023d) and Highmore (2024) focused on the theoretical interpretation and analysis of ICL, which corresponds to Section 5 in this survey. All the above-mentioned survey papers differ with ours in terms of scope and topics. This survey focused on the general development of ICL, including the formal definition of ICL, training strategies, prompt designing strategies, analysis and applications.